### v1

# The DSA Transparency Database: Auditing Self-reported Moderation Actions by Social Media

Abstract
Since September 2023, the Digital Services Act (DSA) obliges large online platforms to submit detailed data on each moderation action they take within the European Union (EU) to the DSA Transparency Database. From its inception, this centralized database has sparked scholarly interest as an unprecedented and potentially unique trove of data on real-world online moderation. Here, we thoroughly analyze all 195.61M records submitted by the eight largest social media platforms in the EU during the first 60 days of the database. Specifically, we conduct a platform-wise comparative study of their: volume of moderation actions, grounds for decision, types of applied restrictions, types of moderated content, timeliness in undertaking and submitting moderation actions, and use of automation. Furthermore, we systematically cross-check the contents of the database with the platforms’ own transparency reports. Our analyses reveal that (i) the platforms adhered only in part to the philosophy and structure of the database, (ii) the structure of the database is partially inadequate for the platforms’ reporting needs, (iii) the platforms exhibited substantial differences in their moderation actions, (iv) a remarkable fraction of the database data is inconsistent, (v) the platform X (formerly Twitter) presents the most inconsistencies. Our findings have far-reaching implications for policymakers and scholars across diverse disciplines. They offer guidance for future regulations that cater to the reporting needs of online platforms in general, but also highlight opportunities to improve and refine the database itself.
Introduction
Content moderation refers to the strategies and processes through which an online platform ensures that the content and behaviors it hosts adhere to the platform’s self-defined standards and guidelines, and to legal regulations (Gillespie 2018). Moderation interventions are the specific actions through which platforms enforce such strategies and processes. Currently, a wide array of interventions are adopted, which involve removing content and users (Jhaver et al. 2021), demoting, restricting, or otherwise reducing their visibility (Le Merrer, Morgan, and Trédan 2021; Trujillo and Cresci 2022), showing users warning messages (Katsaros, Yang, and Fratamico 2022) or attaching warning labels to problematic content (Zannettou 2021), and even acting on monetary rewards (Ma and Kou 2022).
However, despite their importance in maintaining safe and healthy online environments, the application of these interventions often sparks concerns. The controversies are multifaceted, such as considering the potential misuse of removals and reduced visibility decisions as a form of censorship that threatens users’ freedom of speech (Myers West 2018). Furthermore, transparency in content moderation is often questioned, as the processes behind moderation decisions —whether executed by black-box AI algorithms or human moderators— are sometimes perceived as opaque, or lacking clarity and consistency (Gillespie 2020). Issues of explainability, fairness, and the susceptibility of automated systems to biases further contribute to the ongoing debate, highlighting the delicate balance that platforms must strike between upholding community standards and legal obligations, and respecting user rights in the digital sphere (Grimmelmann 2015).
The Digital Services Act. In light of the predominant role held by large online platforms in controlling the flow of information and shaping user opinions, the European Union (EU) recently enacted a strategic response to address the intricate challenges of regulating online spaces. This response also meets the growing demand for legislative intervention to reform the existing model centered on platform self-regulation, wherein social media have long been largely autonomous in defining and implementing their own rules and procedures. The Digital Services Act (DSA) is the regulation officially introduced by the EU on October 2022 to modernize its legal framework governing digital services (European Parliament and Council 2022). Its aim is to create a safer and more accountable online environment for users within the EU, by addressing a range of issues related to online platforms, including content moderation, user safety, and the responsibilities of digital service providers (Turillazzi et al. 2023; Church and Pehlivan 2023). The DSA specifically includes several articles with important implications for online content moderation, the most relevant of which are briefly presented in the following.
  • Article 33 designates those platforms with more than 45M users in the EU as either very large online platforms (VLOPs) or very large online search engines (VLOSEs). Due to their strategic role, these must comply with the most stringent rules of the DSA. On April 2023, the EU designated 17 VLOPs and 2 VLOSEs. [^1] https://ec.europa.eu/commission/presscorner/detail/en/IP˙23˙2413
  • Article 15 demands that VLOPs and VLOSEs release periodic transparency reports. These must contain updated aggregated figures on active recipients, [^2] Those users who either requested hosting of information or were exposed to information hosted on the platform. content removals, user appeals, use of automation and AI in moderation, and timeliness of interventions, among other information.
  • Article 17 obliges VLOPs and VLOSEs to submit a clear and specific statement of reasons (SoR) for each of their moderation interventions. Each SoR must provide detailed and timely information on the intervention, its legal ground, and the content to which it was applied.
  • Article 24(5) establishes that the European Commission must set up and maintain a database of SoRs. On September 2023 the EU established the DSA Transparency Database: an open, standardized, and centralized repository of all SoRs submitted within the scope of the DSA.
Collectively, these and all other DSA articles intend to set the stage for a transparent and harmonized approach to content moderation, underscoring the responsibilities of online platforms while ensuring consistency and fairness. The establishment of the DSA Transparency Database (henceforth DSA-TDB), in particular, marks an important milestone toward enforcing accountability, as it represents an unprecedented resource of self-reported data that enables for the first time the possibility to track, scrutinize, and compare real-world platform moderation actions.
Research Questions and Contributions. Herein, we carry out an in depth quantitative analysis of very large social media moderation practices, as reflected by their SoR submissions to the DSA-TDB. We analyze all 195.61M SoRs submitted by eight platforms during the first 60 days of the DSA-TDB, providing an early yet thorough look into (i) the overall frequency of moderation interventions by very large social media, (ii) their grounds for decision, (iii) the types of applied restrictions, (iv) the types of moderated content, (v) the timeliness of their interventions, and (vi) their use of automation in content moderation. In carrying out these analyses, we specifically focus on answering to the following research questions.
RQ1:
To what extent do social media provide adequate and detailed information in the SoRs they submit?
Article 17 of the DSA specifies that each submitted SoR should be clear, specific, detailed, and timely. RQ1 evaluates the extent to which the different platforms aligned with these criteria. Then, other than shedding light on the adequacy and informativeness of the submitted SoRs, the analysis also uncovered some inherent limitations in the structure of the DSA-TDB that hinder platforms from fully meeting the goals of the DSA.
RQ2:
How consistent and coherent are the SoRs submitted by each social media?
Articles 15 and 24(5) of the DSA establish two transparency mechanisms: transparency reports and the DSA-TDB. By cross-checking information submitted by each platform to the DSA-TDB and by comparing such information with those contained in their DSA Transparency Reports, the analyses entailed by RQ2 reveal the degree to which each platform submitted consistent and coherent information. In particular, we found striking inconsistencies in the information submitted by some of the platforms.
RQ3:
How do social media differ in their self-reported moderation actions?
Assuming that the self-reported information contained in the DSA-TDB are reliable, their analysis sheds lights on the moderation practices of the different platforms. Our results to RQ3 reveal common patterns and marked differences between platforms, along multiple fundamental dimensions of content moderation, including the timeliness of interventions and the use of automation.
Significance. To the best of our knowledge, this is the first platform-wise comparative analysis of moderation actions in a vast social media ecosystem at a continental scale. As such, the implications of our work are many and various. On the one hand, our results allow drawing insights into the moderation practices of different platforms, for example, by identifying common patterns and striking differences. We also uncover incoherent data submitted by some platforms, which raises concerns on the reliability of some of the information contained in the DSA-TDB. On the other hand, our analysis also sheds light on some of the challenges implied by imposing a unified reporting structure to a large number of diverse platforms. Our findings can thus inform future developments of the DSA-TDB, so as to possibly overcome the limitations of the current structure.
Background
The DSA marks a crucial advancement in EU digital services regulation, aiming to enhance online content moderation and citizen protection. It focuses on improving online safety, transparency, and provider accountability (Wilman 2022; Cauffman and Goanta 2021). By emphasizing user protection from illegal content and ensuring fundamental rights, the DSA enforces stringent measures on strategic platforms within the EU —particularly VLOPs and VLOSEs— to reduce risks and foster a safer online space (Husovec 2024).
Recent literature in the realm of digital services regulation offers insightful perspectives on the DSA and on its implications for online content moderation. For example, Urman and Makhortykh (2023) carried out a detailed analysis of the transparency reports released by large online platforms pursuant to Article 15 of the DSA. They found partial compliance with the Santa Clara Principles 2.0 (Santa Clara Principles 2021) among tech companies, with Snapchat fully compliant and Google excelling in state-involved reporting. Their findings also highlighted widespread non-compliance in flagging processes, underscoring a need for enhanced transparency under the DSA. The works by Nunziato (2023) and Church and Pehlivan (2023) explored the global impact that the DSA had on social media content moderation. By comparing the EU’s approach with the DSA to the US’s emphasis on free speech, they suggest that the stringent requirements of the former could set new global internet content moderation norms. Schneider and Rizoiu (2023) went one step further and assessed the potential for harm reduction caused by the DSA, by means of self-exciting point processes. Their analysis, based on real-world data, indicates the DSA’s capacity to standardize and effectively regulate content moderation, with an emphasis on rapid action against highly harmful content.
Instead of focusing on the broad implications of the DSA as a whole, other scholars investigated its possible effects in more specific contexts. For example, Leerssen (2023) investigated the lawful basis for inherently opaque moderation interventions such as demotions and “shadow banning” (Le Merrer, Morgan, and Trédan 2021), revealing a strong tension between platforms’ moderation needs and the transparency-enhancing framework of the DSA. Finally, Strowel and De Meyere (2023) discussed the implications of the DSA in combating online disinformation. The analysis underscored platform transparency requirements and their potential in curbing misleading content. It also contemplated the challenges and effectiveness of these obligations, proposing enhancements to the regulatory framework of the DSA for improved enforcement and compliance.
Overview of the DSA Transparency Database
The main entity in the DSA-TDB is the statement of reasons (SoR) —specific information regarding the removal or restriction to a certain content by providers of hosting services. SoRs are intended to empower users to understand, and challenge, content moderation decisions. Herein, we analyze all SoRs submitted during the initial 60 days since the establishment of the database, from 25 September to 23 November 2023. The data, documentation, and source code of the DSA-TDB are publicly available under permissive licenses on its official website. [^3] https://transparency.dsa.ec.europa.eu/ For our analyses, we used the “light” version of the daily archives, released as compressed CSV files with 33 fields. With respect to the the full database schema, the “light” version excludes four lengthy fields: three free text attributes and one territorial scope attribute. Based on a brief analysis of these four attributes on a sample of a few days of the “full” archive, we decided for the “light” version, as we considered them uninformative in answering our research questions.
Out of the 17 VLOPs designated by the DSA, 16 have submitted at least one SoR. Wikipedia is the only VLOP that did not make any submission to the DSA-TDB during the initial 60-day period. In this time, a total of 551.06M SoRs were submitted. The largest submitter was by far Google Shopping (48.3%). Given our focus on social media —that is, platforms centered on user generated content and social networking— we selected the following eight platforms of interest for this study: Facebook, Instagram, LinkedIn, Pinterest, Snapchat, TikTok, X (formerly Twitter), and YouTube. With 195.61M SoRs, these platforms represent 35.5% of the whole database.
In addition to the DSA-TDB, we also took into account data available on the most recent DSA Transparency Reports released by the eight selected social media (see Table 3 in the Appendix). These compulsory reports mainly cover the first semester of the year (H1 2023) and were all made publicly available around October on the websites of the respective platforms. In particular, we manually extracted the average monthly authenticated active recipients (MAAR) of each platform for H1 2023. MAAR are registered and logged-in active recipients of the service. In the context of moderation, the number of MAAR is more relevant than that of all active recipients, because an active recipient that is not authenticated is usually not allowed to create or generate content on a platform, which is in turn the main target of moderation decisions. Moreover, in the absence of detailed data on the total content created for this period, the average MAAR is a useful, if faulty, metric to relativize the number of SoRs by platform. Table 1 presents an overview of these metrics for the eight selected social media.
[…]
Table 1: Overview of very large social media platforms, ordered by the number of statements of reasons (SoRs) per average monthly authenticated active recipients (MAAR).
Analyses and Results
Upon exploration of the database, we realized that platforms had varying approaches in how they filled in and mapped the attributes of their SoR submissions, especially in terms of mandatory and optional fields. For instance, only 3 out of 16 VLOPs indicate SoR content language —an optional attribute— with none of the social media doing so. Therefore, in this section we present only the aspects which we considered more pertinent in answering our research questions, namely: grounds for the decision on an infringement, types of restriction decisions, type of moderated content, timeliness of intervention application and communication to the DSA-TDB, and the use of automated means for moderation. As a final validation step, we compare the data submitted by each social media to the DSA-TDB with the respective claims from their DSA Transparency Reports, so as to detect possible inconsistencies.
Grounds for Decision
There are several attributes to indicate the legal or contractual grounds relied on in taking a moderation intervention decision, and to codify the infringement type. We first considered two grounds attributes: a mandatory decision ground (illegal, incompatible), and an optional attribute to indicate that a content is both incompatible with the terms of service and illegal. We then analyzed three attributes to codify infringements: category (mandatory high-level attribute from a predefined set of 14 categories), category addition (optional list from the same set of categories), and category specification (optional keyword list from a set of 55 predefined values).
Concerning the decisions grounds, for almost every platform incompatible content was the overwhelming majority, up to 100% in the case of LinkedIn. The exception was X, which indicated illegal content for all of its SoRs. Only Snapchat indicated an overlap of incompatible and illegal content in 12.2% of their SoRs. None of the other social media platforms used this attribute.
Figure 1: Platform-wise distribution of infringement categories of statement of reasons (SoR), seriated using Manhattan distance across platforms and categories, so that similar platforms are positioned close to one another.
Figure 1 provides a comparative analysis of the platforms based on the infringement categories of their SoRs. As shown, there exist three groups of similar platforms:
  1. 1.
    those in which scope of platform service [^4] This is a generic category to indicate content moderated because deemed outside of the scope of the platform service. was the most common infringement (YouTube, Facebook, LinkedIn, TikTok, and Instagram);
  2. 2.
    those in which pornography or sexualized content was the most common infringement (Pinterest and X);
  3. 3.
    a single-platform group (Snapchat), for which the previous categories are the most common, along with scams and fraud.
Curiously, category addition was merely used by Pinterest and X, but in less than 0.04% of their SoRs. As for keywords, only half of the social media platforms used them for category specification: Pinterest, Snapchat, X, and YouTube. However, these are not always informative. For instance, circa 75% of YouTube’s SoRs have a keyword —all of them other— which YouTube explained by specifying that “[their] mapping has not yet been completed”. Only 10% of Snapchat’s SoRs have a keyword, 86% of these being sexual abuse material and the rest online bullying intimidation. About 87% of Pinterest’s SoRs have a keyword (with a few having two), predominantly adult sexual material. Interestingly, all of X’s SoRs have a keyword (with a few having two or three), but almost half (45%) is other (detailed as a policy violation), while 38% is adult sexual material. The rest is child sexual abuse material, nudity, and goods/services not permitted.
Decision Type
visibility account provision monetary
[…]
Table 2: Platform-wise distribution of decision types, ordered alphabetically and by type mean across platforms.
Figure 2: Platform-wise distribution for decisions of the visibility type, seriated using Manhattan distance across platforms and sub-types, so that similar platforms are positioned close to one another.
Figure 3: Distribution of moderated content type by platform. Custom other names were simplified for space reasons.
Figure 4: Distribution of decision application delays (with respect to the date of content creation) by social media platform, ordered by their 80% winsorized means (W).
Types of Restriction Decisions
There are four main types of restriction decisions: account (suspension, termination), monetary (suspension, termination, other), provision (partial/total suspension/termination of the service), and visibility (restriction for the content). Platforms must indicate a single sub-type of one of the four types. However, these types are not mutually exclusive, meaning that a SoR can contain multiple types of restrictions, with at least one of them being mandatory. However, all of the social media platforms set a single type. As detailed in Table 2, most were in visibility (78.7% across platforms), followed by account (18.1%) and provision (3.1%). Surprisingly, there was not a single SoR with a monetary restriction decision. For visibility sub-types, shown in Figure 2, we noticed a similar platform grouping to that of infringement categories:
  1. 1.
    those in which content removed is the most common sub-type (Facebook, Instagram, LinkedIn, TikTok, and YouTube);
  2. 2.
    those in which other is the most common sub-type (X and Pinterest);
  3. 3.
    Snapchat by itself, with content disabled being its most common sub-type.
Content Type
The required attribute content type is a list with one or more of the following predefined values: app, audio, image, product, synthetic media (e.g., content generated via AI), text, video, and other. Despite the possibility to indicate multiple values, all eight platforms only specified a single value for all of their SoRs, except YouTube, which however indicated multiple values only in a handful of cases (<<<0.01%). We found that social media platforms greatly differ in how they specify the content type of their SoRs, as shown on the left side of Figure 3. Pinterest and LinkedIn only set content type other, and did not map any of their internal types to one or more of the predefined values. YouTube also had a remarkable share of other content (91.1%), prevalently ads. Facebook and Instagram have similar distributions, with most content being other (almost 60%), all detailed as accounts, which is consistent with their distributions of decision types in Table 2. Incidentally, Facebook was the only social media with content type product (1.2%). Most of Snapchat’s moderated content was video, with other content (24%) being mainly accounts. Still, 16.6% of other was labelled as multi-media, instead of using multiple predefined values as prescribed by the DSA-TDB. Interestingly, TikTok’s moderated content was mostly text (55.5%) and not video (33.8%), with a small share of other content (1.3%), primarily accounts. Unexpectedly, moderated content by X was overwhelmingly synthetic media (99.8%), with the second being audio (0.01%) —the only social media to use this type— and only a handful of text.
Timeliness
Artice 17 of the DSA requires that platforms communicate their SoRs without undue delay to allow close to real-time updates. Pinterest, Snapchat, and TikTok started submitting SoRs on the very day that the DSA-TDB was made available (i.e., on 25 September). Facebook, Instagram, YouTube, and LinkedIn the following day. X was last in line and started a week later (i.e., on 3 October).
Figure 5: Platform-wise distribution of automated infringement detection and decision, ordered by an ad hoc automation index.
We then analyzed the delay with which social media communicated SoRs, measured as the difference in days between the date of application of the moderation intervention and the date of communication in the corresponding SoR (see Figure 8 in the Appendix). LinkedIn, Pinterest, and Snapchat for the most part communicated their SoRs on the same day of application and with little to no fluctuations. Despite the late start, X caught up on the first week of submissions by communicating content moderated since 25 September. Afterwards, it began to communicate SoRs the same day of their application. On the contrary, we noticed that TikTok started by stably submitting on the same day of application, but at the end of the 60-day period it communicated SoRs with an uprising mean delay of up to six days. On the other hand, Facebook, Instagram, and YouTube communicated their SoRs with a mean delay of one day, but with marked fluctuations throughout the period.
Arguably, a prompt response by platforms in moderating an infringing content is desirable. Hence, we also measured the delay in days of the restriction decision with respect to the creation of the infringing content. We observed great heterogeneity in the distributions of these delays among platforms. For this reason, in their analysis and visualization we opted for using few categorical cuts, as shown in Figure 4. In addition, we also report in the figure the 80% winsorized mean (W) as a robust measure of centrality that lessens the influence of outliers. Oddly, X was the only platform that always applied a decision on the same day the infringing content was created (W===0). TikTok had the second lowest delay, with 89.7% of decisions being applied the day of content creation. Next, LinkedIn applied its decision within 7 days in 89.6% of their SoRs, while Instagram in 81.4%. Even if Snapchat had a higher value for the same timeframe (82.5%), in 14.1% of SoRs it took more than 30 days to apply its decision. Facebook presented a similar but even more marked situation, with 84.2% and 22.5%, respectively. YouTube, on the other hand, had a substantial amount of decisions that took more than 30 days (68.4%), while also having many within 15–30 days (15.3%). However, the distribution with the highest mean and dispersion was that of Pinterest, as 16.9% of decisions were applied the same day of content creation while 71.6% after 30 days, with many even taking hundreds of days.
[…]
Automated Means
Automation can be applied at various steps in the content moderation process. It is often used to scale up the detection of offending content and to speed up the decision on how to moderate it. Consequently, the DSA-TDB includes two attributes to describe the possible use of automation for moderation: automated detection (yes, no) and automated decision (not, partially, fully). Figure 5 provides a combined view of the use of automation for detection and decision by social media platforms. Again, we see different approaches being adopted. Surprisingly, none of the SoRs submitted by X specified the use of automated means whatsoever. Snapchat is second in terms of non-automated means in their SoRs (76%). Pinterest mainly uses non-automated detection with partially automated decision (98%), whereas LinkedIn mainly uses automated detection together with non-automated decision (68%). YouTube is the only platform that uses all combinations of automated means, with automated detection and non-automated decision being most common (38%), followed by fully automated means (31%), and non-automated means (28%). Once more, Instagram and Facebook have similar distributions, with automated detection and partially automated decision being the most common for both, with 92% and 99%, respectively. Almost opposite of X, TikTok predominantly indicates fully automated means for moderation (92%).
To provide a holistic view of social media behavior with respect to automated moderation, we converted the boolean and ordinal attributes of the DSA-TDB into two numeric indices of automated detection and automated decision. Both indices are defined in the [0, 1] range and are valued, for each platform, depending on the frequency with which automated detection and decision occurred in the platform’s SoRs. We were then able to position each platform in the bi-dimensional space depicted in Figure 6, where dot size is proportional to the number of SoRs per MAAR, and dot color represents the mean of the automated detection and automated decision indices. The platform positioning within the newly defined automation space reveals four groups of platforms with overall similar degrees of automation: low (X and Snapchat), modest (Pinterest), substantial (LinkedIn, YouTube, Instagram, and Facebook), and high (TikTok).
Comparison with Transparency Reports
Within the DSA framework there are multiple mechanisms to ensure transparency in content moderation, including the DSA-TDB that we analyzed so far, and the DSA Transparency Reports. Both resources cover similar aspects of the content moderation process, albeit at different granularity. As such, there is overlapping between the information contained in the two resources, which opens up the possibility to carry out thorough comparisons. In this section we leverage this opportunity to carry out the first systematic comparison between the data submitted by each social media to the DSA-TDB and that declared in the corresponding transparency report. In detail, for each platform and for each dimension of moderation (e.g., timeliness, automation, etc.), we look for the corresponding information in that platform’s most recent transparency report. We then compare the two figures to corroborate or challenge the self-declared information provided by each platform. This comparison serves as a validation step of the DSA-TDB data, and allows for spotting inconsistencies and discrepancies. In general, we were not able to compare each dimension of moderation for each of the eight platforms, due to differences in the data they provided in their transparency reports. However, all platforms —with the exception of Snapchat— provided detailed information about the use of automated means for moderation, for which we found many inconsistencies. In the case of X, however, we found inconsistencies in multiple aspects. In the following, we present the results of our comparisons starting from the platforms for which we found less issues.
Pinterest, Facebook, Instagram, LinkedIn. Pinterest’s self-reporting was overall congruous and detailed. Interestingly, it was the only social media to use three automation levels in their transparency report —manual, hybrid, and automated— similarly to the organization of the DSA-TDB. Unsurprisingly, Facebook and Instagram share similar transparency reports, based on Meta guidelines. In their reports, both state the heavy use of automated means, including full automation for both for detection and decision. However, none of their SoRs specified fully automated decisions, let alone full automation overall, as shown in Figure 5. For the reporting period 25 April – 30 September (159 days), Meta platforms automatically removed the following organic content (i.e., non-paid): for Facebook, 46.68M pieces of content (93.94% of the total), and for Instagram 75.11M (98.44% of the total). These proportions of automatically removed content are consistent with their SoR submissions only if there is no distinction between fully and partially automated decision. LinkedIn’s transparency report covers moderation actions occurred between 25 August and 30 September. It distinguishes between job posts and other content, with the former having a more stringent moderation policy. In its report, LinkedIn claimed to have removed all kinds of violating content detected automatically 99% of the time, whereas for submitted SoRs it was 86% —a relatively small discrepancy.
TikTok, YouTube, Snapchat. TikTok’s transparency report lacks several details and their actions data only covers the month of September 2023, shortcomings that are acknowledged therein. Based on the more detailed actions for removed content, we found a considerable automation discrepancy. The report specifies that 45% of non-ad content was removed automatically, whereas in the DSA-TDB it was 95%, more than double for a partially overlapping period. In the case of YouTube, even if it is the only platform to use all combinations of automated means without one being predominant —in accordance with its transparency report— we found an intriguing anomaly during our exploratory analysis. In the DSA-TDB data we found a remarkable spike in SoRs applied on a single day —5 October— reaching 732.4K. For reference, YouTube’s daily median is 112.2K. Almost all of these are ads that were disabled because deemed outside the scope of the platform, but via non-automated means, which seems implausible given the very large amount of moderation actions undertaken. Unfortunately, with the SoR and transparency report data available we could not delve into this anomaly in more detail. The DSA transparency report by Snapchat is somewhat peculiar, as they already had in place a global transparency report, released twice a year. Hence, they only integrated this report via a document with some high-level information specific to the EU, in order to comply with the first mandatory DSA reporting period. Nevertheless, we noticed an odd inconsistency. DSA transparency reports must include data about Trusted Flaggers: legal entities responsible for notifying platforms of illegal content. As per Article 19(2) of the DSA, Trusted Flaggers are officially appointed by certain EU bodies. However, at the time of writing no Trusted Flagger has yet been designated. For this reason, no platform reported in their SoRs to have received notifications from Trusted Flaggers. With the exception of Snapchat, for which we found a few hundreds SoRs with notification source type set as Trusted Flagger. Upon further inspection we realized that Snapchat has a similar namesake internal program, [^5] https://values.snap.com/safety/safety-center unrelated to the DSA’s. This homonymy probably caused some confusion in Snapchat’s reporting to the DSA-TDB, where they referred to their internal program instead of the intended one.
X. Finally, the transparency report by X includes very detailed data regarding its moderation actions for the period 28 August – 20 October (54 days), with 56K actions being taken on content (36.5% via automated means) and 2M actions on accounts (74.3% via automated means). Yet, all of their SoRs submitted to the DSA-TDB are marked as not using any automated means. In addition, during the 60-day period of our study —which partially overlaps the transparency report’s period— X only submitted 264K SoRs, compared to 2.05M moderation actions declared in the transparency report. These actions covered 14 violation categories, which are similar but not the same as the DSA-TDB categories. For content, violent speech and abuse & harrassment were the most common: 25.5K (47%) and 12.1K (22%), respectively. For accounts, platform manipulation & spam and child sexual exploitation: 1.89M (95%) and 60K (3%) violations, respectively. However, 99.8% of all SoRs submitted by X to the DSA-TDB targeted synthetic media, mostly pornography or sexualized content, as shown in Figures 3 and 1, which appears to be inconsistent with X’s transparency report. Moreover, all of X’s SoRs in the DSA-TDB reported a decision application delay of zero days, without any use of automation, which again seems implausible.
Discussion
We first discuss the adequacy of the data submitted to the DSA-TDB (RQ1). We then examine inconsistencies in the submitted data (RQ2) and we describe interesting differences and similarities among platforms (RQ3). Finally, we ponder on the potential misuse of the database contents in its current state, for research purposes and policymaking.
RQ1: Adequacy of Submitted Data
We noticed several inadequacies in the data of social media platforms. In several cases the attributes in the database were conceived to allow platforms a higher level of granularity in characterizing a SoR, compared to what is actually done. For instance, there is a severe lack of detail regarding the types of infringements. The optional category addition is practically never used, and the rich set of available keywords in category specification is severely underutilized. Further, in most cases platforms specified the infringement category as the least informative scope of platform service. Similarly, platforms have the possibility to apply more than one of the four types of restriction decisions to a SoR, but none of the social media did so. In particular, it was surprising to see that no monetary restriction decision was specified whatsoever, especially in platforms with well-established content monetization schemes subject to strong moderation, such as YouTube (Ma and Kou 2022). Likewise, content type allows multiple values, but in the great majority of instances only one was set. This is particularly problematic with synthetic media, as it would be important to also know the kind of media of the generated content, whether text, audio, video, or a combination of them (Alam et al. 2022). These widespread deficiencies not only diminish the utility of the DSA-TDB by limiting the depth of available data, but also compromise its commitment to transparency.
In general, the initial implementation of the SoR communication systems by social media platforms was not as comprehensive as it could and should be. Content language is an illustrating example. The language in which moderators work is one the key aspects that VLOPs must cover in their transparency reports —an expected requirement given the multilingual policy to which the EU strives (Gazzola 2016). Still, none of the social media platforms set the content language in their SoRs. From a technical point of view most, if not all, of these very large social media already have mechanisms to register or automatically detect the language of a piece of content or account (e.g., via the user interface settings). Moreover, this attribute is essential for the platform, not only for the the moderation via manual our automated means (Lees et al. 2022), but also to sell targeted ads and generate revenue (Beauvisage et al. 2023). We presume platforms avoid communicating this attribute —and the resulting workload— simply because it is optional.
That being said, we believe that some fault also lies in the shortcomings of the current database schema. We acknowledge the challenges entailed in the definition of a common and unifying schema, not only for several social media, but for all the kinds of online platforms covered by the DSA. In fact, in mid 2023 there was a public consultation on the DSA-TDB, in which 48 stakeholders participated, including representatives of 6 VLOPs. The consultation was intended to refine an existing draft of the schema and SoR communication mechanisms. This also means that VLOPs only had a few months to implement these changes. Nonetheless, we noticed that certain attribute definitions could be improved, based on their actual use. For instance, many platforms apply their restriction decisions at the account level, not because of a specific type of multi-media content created by it, but due to the account itself. However, in the initial database schema it was not possible to specify an account as the target of a restriction. Hence, many platforms indicated the SoR content type as other and detailed it as an account-based type, as we showed in Figure 3. Perhaps, for the future it would be better to separate the target kind at least into two attributes: content and account, both of which could then be further specified, as already occurs for content type and the many account variants in other. Appropriately enough, the EU has undertaken an open approach to further develop the DSA-TDB and states it is willing to receive and act upon feedback.
RQ2: Data Inconsistencies
We found several inconsistencies in the self-reporting of social media platforms, based on a comparison between their DSA Transparency Reports and their SoRs. According to our comparative analysis, Pinterest resulted to be the most consistent in this regard, while also offering detailed data on its moderation actions. YouTube was also consistent in its self-reporting data, albeit the prominent spike of restricted content on 5 October casts doubts on the completeness of the submitted data. The peculiar inconsistency of Snapchat, which indicated hundreds of notifications of illegal content by Trusted Flaggers when none has been yet designated under the DSA, seems to be an honest mistake due its internal namesake program. Nonetheless this peculiarity stands as an example of the complexity that emerges when adapting already established internal processes to new and vast norms.
The majority of inconsistencies that we found was related to the use of automated means. In some cases, such inconsistencies appear to be a matter of interpretation. For instance, Facebook states in its transparency report: “[e]very day, we remove millions of violating pieces of content and accounts on Facebook. In most cases, this happens automatically, with technology to detect, restrict, and remove content and accounts […]”. Later, they further specify: “when confident enough that a post violates one of our Community Standards, the artificial intelligence will typically remove the content or demote it. We also use artificial intelligence to select the content for human review on the basis of severity, virality, and likelihood of a violation”. Instagram’s report includes similar statements. In other words, both Meta platforms declared heavy use of fully automated means in their transparency reports, but not a single fully automated action is reported in their submissions to the DSA-TDB, as is evident in Figure 5. Instead, they opted to indicate partially automated detection by an overwhelming majority. Perhaps this choice was based on Meta’s declarations of “us[ing] human reviewers to assess the accuracy [of automated means] against current content, behaviour, or accounts, rather than just historical ones”. Regardless of the reason, we beg to differ, as we believe that at least some SoRs should have been classified as fully automated, because —as implied by Meta itself— there is no human involvement during the moderation action itself.
In other cases, the inconsistencies are a mismatch between the reported share of automated actions between the platforms’ transparency reports and the DSA-TDB. Some inconsistencies are relatively small, as in the case of LinkedIn whose reported use of full automation differ by circa 10 percentage points. Other times the inconsistencies are much larger, as for TikTok, whose differences are in the region of 50 percentage points.
Instead, X emerges as a distinctive outlier and as the most prominent case of self-reported inconsistencies. In its transparency report, the platform states: “X employs a combination of heuristics and machine learning algorithms to automatically detect content that violates the X Rules and policies […]”. Later they reiterate: “[c]ontent flagged by these machine learning models are either reviewed by human content reviewers before an action is taken or, in some cases, automatically actioned based on model output”. Yet, all of the SoR submissions by X during the period of study specified non-automated detection and decision. Not even a single partially automated decision was reported, as shown in Figure 5, which represents a striking inconsistency. In addition to the lack of use of automated means, other discrepancies include a lower than expected number of moderation actions, the unlikely predominance of illegal content, as well as a single value for both content type and restriction decision. These discrepancies in number and characterization cast serious doubts on the reliability of the SoRs submitted by X. On the contrary, its transparency report is arguably one of the most, if not the most, comprehensive of the eight platforms, with detailed tables and explanations regarding their moderation actions.
Figure 7: Bi-dimensional Sammon mapping of very large social media platforms, based on the Manhattan distance of ten of their most relevant self-reported moderation metrics. The green diamond shape marks the platforms’ centroid.
RQ3: Platform Differences
As already mentioned, we found considerable differences in the way the platforms specify their SoR submissions to the DSA-TDB. For illustrative purposes, in Figure 7 we used ten metrics from the presented analyses to visualize the similarity among very large social platforms in terms of their MAAR and SoR submissions. We utilized Sammon mapping, a non-linear multidimensional scaling technique commonly used in exploratory data analysis (Saeed et al. 2018), setting the Manhattan distance as a similarity measure.
From this visualization, the dissimilarity of X with respect to the other platforms is evident. Indeed, we found that X greatly differed from the other platforms in terms of automated means, decision grounds, as well as content and restriction types and sub-types. However, we also found that X was the platform with more inconsistencies with respect to its DSA Transparency Report. Pinterest is the second largest dissimilar platform, mainly due to their remarkable high average and dispersion in the delay between content creation and application of the moderation, as shown in Figure 4. Another difference lies in its high share of actioned pornography or sexualized content. TikTok, somewhat diverges due its high volume in SoRs (52.69% of all social media platforms), SoRs per MAAR, and predominant use of fully automated means, as reported in Table 1 and Figure 5. Whereas LinkedIn does so because it has the lowest volume of SoRs, MAAR, and SoR per MAAR. Unsurprisingly, Facebook and Instagram are the closest to each other. According to their transparency reports, they apply the same or very similar moderation policies, in line with Meta guidelines, which is reflected in their SoRs. Their main dissimilarity lies in the volume of moderation actions, with Facebook having almost ten times more SoRs for a similar number of MAAR than Instagram. Finally, Snapchat and YouTube share many characteristics with the other platforms, given the more balanced specification of their SoRs.
Differences among social media in terms of SoRs are not only the result of their moderation policies, but also, and perhaps foremost, due to the scope of the platforms themselves. Different social media cater for different users by means of their user experience, content offering, and community dynamics. Consequently, social media users usually browse more than one platform, each at distinct moments, for different durations, and with diverse motivations (Alhabash and Ma 2017). These intrinsic differences should be taken into consideration when analyzing the content of the DSA-TDB. For instance, in our exploratory analysis we measured the seasonality strength of the number of daily interventions on a scale from 0 to 1, based on iterative STL (i.e., seasonal and trend decomposition using LOESS). LinkedIn had a very strong weekly seasonality of 0.81 on a 0 to 1 scale, almost double of the second one. In fact, we found a marked decline in LinkedIn SoRs with application date on weekends, while peaks usually occurred mid week. This is likely due to two factors: (i) there is less human moderator activity on weekends —which is relevant considering that only 17% of LinkedIn decisions are automated— and (ii) users mostly use the platform on weekdays while at work —given LinkedIn’s focus on business and employment. This is but an example and complementary data sources are needed to better understand this phenomenon. However, it serves the purpose of showing the great heterogeneity of platforms and their moderation actions.
Final Remark: A Word of Caution
Great care is needed when conducting research studies based on the initial data of the DSA-TDB. Above all, one must not forget that these data are self-reported. Even if SoR submission and transparency reporting are mandatory under the DSA, so far the reliability of the data lies on the willingness and the resources of the platforms to be comprehensive and accurate in their mapping of existing moderation processes to the novel and complex norms. As we have shown, this is not yet the case for the nascent DSA-TDB. Indeed, our analyses indicate that social media SoRs are not adequate with respect to the objectives of the database, manifest critical inconsistencies, and reflect platforms with very different scopes, domains, and dynamics. Consequently, there is a high risk of making misleading inferences based on these data if the appropriate precautions are not taken, such as using complementary data sources for validation. Moreover, it is possible that these issues will persist or even be exacerbated once all providers of online platforms subject to the DSA —and not only VLOPs and VLOSEs— will be obliged to submit their moderation data to the DSA-TDB. If VLOPs have these issues, we expect that smaller and less resourceful platforms will struggle to do better.
Conclusions
The recently established DSA Transparency Database is a trove of data on the moderation actions of large social media platforms, allowing the unprecedented opportunity to audit and compare their moderation practices. Here, we analyzed all 195.61M statements of reasons (SoRs) submitted by eight social media platforms during the first 60 days of the database. Other than providing one of the first accounts of the content of the database, our analyses also compared the reported moderation actions of the different platforms with what they declared in their transparency reports. We uncovered striking differences, inadequacies, and inconsistencies in the data submitted. Based on our results, we conclude that large social media adhered only in part to the philosophy and structure of the database. Specifically, while all platforms met the requirements of the DSA, most omitted important (yet optional) details from their SoRs. This limits the usefulness of the database and hinders further analysis of this data by scholars and policymakers alike. At the same time, the structure of the database turned out to be partially inadequate of the platforms’ reporting needs. For example, the current structure lacks the possibility to explicitly report moderation decisions targeted at accounts rather than content. To this regard, our results can inform future developments of the database.
Social media platforms exhibited marked differences in their moderation actions, in terms of restriction and content types, timeliness, and automation. Part of the differences are due to the platforms’ varying degrees of adherence to the database. However, part are also due to the unique characteristics of each platform, which pose challenges when it comes to harmonizing their moderation practices, as envisioned by the DSA. Furthermore, we also found that a significant fraction of the initial database data is inconsistent. This resulted from both internal validity checks on the data itself, as well as from external checks against statements and data extracted from the platforms’ transparency reports. Of all the considered social media, X presents the most inconsistencies, as reflected by striking discrepancies found in multiple aspects of its moderation actions. Finally, we conclude that the self-reported nature of the database, and the widespread inconsistencies that we found, raise concerns on the reliability of the data contained therein. This begs caution when analyzing such data to make inferences and draw conclusions on the moderation practices of large social media platforms.
Appendix
Sources of DSA Transparency Reports
Table 3 contains the URLs to the documents or repositories of the DSA Transparency Reports for the analyzed social media platforms. Beware that platforms differ in the periods of reference and publication dates in their reports. We used those that mainly focus on H1 2023 and were published around October of the same year. In some cases the active recipients of a given platform are detailed in a different document within the respective repository.
Delay in Communication of Moderation Actions
Figure 8 is the time series of the daily mean delay in communicating SoRs to the DSA-TDB. The delay (measured in days) is the difference between the creation date of the SoR and the application date of the restriction decision. Of interest are: the near-zero and stable delay of LinkedIn, Pinterest, and Snapchat; the increase in delay for TikTok at the end of the period of study; the way in which X caught up after being one week behind schedule; and the increasing fluctuations of YouTube toward the end.
Daily Count of Moderation Actions
Figure 9 is the time series of the daily count of SoRs based on the application date of the decision for the action. Not too much attention should be paid to the values at the start and end of the time series. For instance, some platforms began their SoR communication a few hours late, which in part explains the seemingly low count at the start. In the case of Facebook and Instagram, they usually communicate their SoRs the day after their application date, hence the apparent sharp decline at the end of the period. The differences in daily volume of SoRs among the social media platforms is noteworthy. In this regard, there is a noticeable upward trend in Facebook, which began with a surge in moderation actions on 7 October, the day of the Hamas attack on Israel. A similar surge can be appreciated on Instagram, but with a stable yet higher trend afterward. Despite their relevance, we did not delve into these phenomena because these are outside the scope of our study. Also of interest are: the strong weekly seasonality of LinkedIn and to a lesser degree that of Pinterest; the fall and rise on TikTok and X; and the YouTube spike on October 5, an anomaly in terms of volume and characterization of SoRs for that day.
Figure 8: Platform-wise daily mean of the delay (measured in days) in the communication of SoRs to the DSA-DTB.
Figure 9: Platform-wise daily count of SoRs by application date of the restriction decision.